# MMD-Sense-Analysis: Word Sense Detection Leveraging Maximum Mean Discrepancy

### v1

@@ 2.1 MMD Two-Sample Variable Selection
Maximum Mean Discrepancy (MMD). Maximum Mean Discrepancy (MMD) is a kernel-based distance metric between probability distributions Gretton-2012a and a measure of distance between two probability distributions. It relies on a positive semi-definite kernel Hofmann-2008 ; Smola-2009 to map the distributions into a new space, a Reproducing Kernel Hilbert Space (RKHS), where probability distributions are represented as data points Muandet-2017 . A large MMD value indicates significant differences between two distributions. I depict an intuitive presentation of the MMD in Figure 1.
Figure 1: An intuitive illustration of the MMD.
For a kernel function $k:\mathbb{R}^{D}\times\mathbb{R}^{D}\to\mathbb{R}$, the MMD between probability distributions $P$ and $Q$ on $\mathbb{R}^{D}$ is defined as
${\rm MMD}_{k}^{2}(P,Q):=\mathbb{E}_{{\bm{x}},{\bm{x}}^{\prime}{\sim}P}[k({\bm{x}},{\bm{x}}^{\prime})]+\mathbb{E}_{{\bm{y}},{\bm{y}}^{\prime}\sim Q}[k({\bm{y}},{\bm{y}}^{\prime})]-2\mathbb{E}_{{\bm{x}}\sim P,\ {\bm{y}}\sim Q}[k({\bm{x}},{\bm{y}})],$
where ${\bm{x}},{\bm{x}}^{\prime}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}P$ and ${\bm{y}},{\bm{y}}^{\prime}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Q$ are random vectors, and the expectation $\mathbb{E}$ is taken for the random vectors in the subscript. This is a proper distance metric on probability distributions, if the kernel $k$ is characteristic (e.g., when $k$ is Gaussian) (Sriperumbudur-2010, ). Given i.i.d. samples ${\bf X}=\{{\bm{x}}_{1},\dots,{\bm{x}}_{n}\}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}P$ and ${\bf Y}=\{{\bm{y}}_{1},\dots,{\bm{y}}_{n}\}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Q$, the MMD can be consistently estimated as
@@
2.2 MMD-Sense-Analysis: Detecting and Analyzing Word Sense Changes
@@ Acknowledgements
I gratefully acknowledge Mr. Hajime Kiyama of Hitotsubashi University for generously providing the word embedding models used in this study.

### v2

# Word Sense Detection Leveraging Maximum Mean Discrepancy

@@ 2.1 MMD Two-Sample Variable Selection
Maximum Mean Discrepancy and MMD Variable Selection.
Maximum Mean Discrepancy (MMD) is a kernel-based distance metric between probability distributions Gretton-2012a and a measure of distance between two probability distributions. It relies on a positive semi-definite kernel Hofmann-2008 ; Smola-2009 to map the distributions into a new space, a Reproducing Kernel Hilbert Space (RKHS), where probability distributions are represented as data points Muandet-2017 . A large MMD value indicates significant differences between two distributions.
Given i.i.d. samples ${\bf X}=\{{\bm{x}}_{1},\dots,{\bm{x}}_{n}\}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}P$ and ${\bf Y}=\{{\bm{y}}_{1},\dots,{\bm{y}}_{n}\}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Q$, the MMD can be consistently estimated as
@@
2.2 Detecting and Analyzing Word Sense Changes
@@ Acknowledgements
This work has been supported by the French government, through the NIMML project (ANR-21-CE23-0005-01). I gratefully acknowledge Mr. Hajime Kiyama of Hitotsubashi University for generously providing the word embedding models used in this study.